OpenStudy (anonymous):

Simplify and determine the coefficient of (-x)(3y)(-2x).

OpenStudy (anonymous):

Look at the coefficients of each first. -1*3*-2=6 Next look at the x's: x*x=x^2 Finally there's just a y All of them together is 6*x^2*y :)

OpenStudy (anonymous):

thanks